Write a slope-intercept equation for a line parallel to the line x – 5y = 9 which passes through the point (10, –3).

Answer by stanbon(75887)

Write a slope-intercept equation for a line parallel to the line x – 5y = 9 which passes through the point (10, –3).
----
5y = x-9
y = (1/5)x - (9/5)
----
slope = 1/5
---
Answer Form:: y = mx + b
Solve for "b" using y = -3, x = 10, m = 1/5
-3 = (1/5)10 + b
b = -5
------
Equation:
y = (1/5)x - 5
------------------
Cheers,
Stan H.
--------------